Two identical, uniform rods are configured as shown above, and the assemble is released at rest (in the position shown). Find, immediately after release:

\alphabc (angular acceleration) (magnitude and direction)

ab (magnitude and direction)

rods 45 Min OX unium and the asbl ured as shown above wo idenicaods are confiq is released at test ántaerosttionshow^1. Find, immediately after elease:
Answer #1

Inertia = I = 1/3 * m*l^2 = 1.8*0.840^2/3 = 0.42 kg m^2

m = mass = 1.8 kg, l = length of the rod = 0.84

Torque = I* alpha

alpha = angular acceleration

Torque = mgx

g = acceleration due to gravity = 9.8 ms^-2

x = perpendicular distance between the line of action and mass

x= l/2 cos (45)= 0.84/2 cos (45) = 0.29 m

Torque = mgx=1.8*9.8*0.29 = 5.11 Nm

alphaBC=angular acceleration of BC = torque/Inertia = 5.11/0.42 = 12.2 rad/s^2 (in the direction perpendicular to the plane of the paper using right hand thumb rule)

aB = linear acceleration = 0 (because the point is not moving)